# Towards Generalized Source Tracing for Codec-Based Deepfake Speech

### v2

@@ I Introduction
Deepfake detection determines whether the given speech is a bona fide speech or a deepfake speech. Numerous challenges and datasets [1, 2] have been introduced, such as ASVspoof [3, 4, 5, 6, 7] and ADD [8, 9]. Recently, attention has shifted from merely detecting deepfake speech to tracing its source. Source tracing offers important insights into the origin or generation algorithm of a deepfake speech, which is especially valuable in out‐of‐domain detection scenarios. For example, Track 3 of ADD 2023 [9] focused explicitly on identifying the underlying algorithms of deepfake utterances, and Interspeech 2025 is hosting a special session [^1] https://deepfake-total.com/sourcetracing about deepfake source tracing.
The simplest approach to source tracing [10] involves mapping deepfake utterances back to the specific attack model ID that generated them [9]. To enhance generalization, the field has shifted toward attribute‐based source tracing [11, 12, 13, 14]. Rather than treating each deepfake system as a unique category, this method groups systems according to shared attributes, reflecting the fact that many deepfake pipelines reuse common components. By focusing on these attributes, models can more effectively trace spoofing algorithms that were unseen during training but are built from known elements during training [14]. Drawing on well‐defined properties from traditional text-to-speech (TTS) and voice conversion (VC) architectures, researchers have concentrated on classifying attributes such as input types [14], acoustic models [12, 14], speaker representations [13], and vocoders [13, 11, 12, 14], etc.
However, these prior studies target source tracing in traditional TTS/VC systems, and only initial efforts have explored codec‐based speech generation (CoSG) [15] models. Xie et al. [16] treat source tracing as a closed‐set classification problem—identifying known CoSG system IDs and using a binary out‐of‐distribution detector, but this approach fails when multiple unseen generators are encountered. Although there has been a series of recent works on source tracing [17, 18, 19], few focus on codec-based deepfake speech [20, 21]. More recently, Chen et al. [22] utilize the organized CodecFake+[21] taxonomy, which covers vector quantization schemes, auxiliary objectives, and decoder types, to formulate three source‐tracing tasks grounded in codec attributes. They [22] also show that a model trained on codec‐resynthesized data suffers from severe generalization issues when evaluated on CoSG‐generated speech, driven by unseen codec types.
@@ II Background
Neural audio codecs [23, 24, 15, 25, 26] leave “fingerprints” across three axes: vector quantization, auxiliary objectives, and decoder type. Together, these axes span most codec designs and enable robust tracing of codec-based deepfake. Building on the taxonomy, Chen et. al. [22] define three multi-class classification tasks for CodecFake source tracing:
  • Vector Quantization Classification (VQ Task): Four-class classification: classifies a waveform as real or generated by a model with codec using multi-codebook, single-codebook, or scalar quantization.
  • Auxiliary Objective Classification (AUX Task): Three-class classification: classifies a waveform as either real or generated by a model with a codec using semantic distillation, or disentanglement objectives.
  • Decoder Type Classification (DEC Task): Three-class classification: classifies an input utterance as real or generated by a model with codec employing a time- or frequency-domain decoder.
Additionally, a traditional anti-spoofing task is included:
  • Binary Spoof Detection (BIN Task): Classifies each input utterance as bona fide or spoofed speech.
@@ III-A Semantic Encoder
The Semantic Encoder leverages a frozen, pre‐trained Whisper [29] model to generate content‐robust embeddings that serve as a stable anchor for source tracing. By verifying that two utterances share identical linguistic content, these embeddings ensure the network ignores natural acoustic variations caused by different words or pronunciations and focuses only on codec‐induced distortions. Concretely, let ${x}$ denote the raw input waveform. We first zero‐pad ${x}$ to a fixed duration of 30 seconds, yielding ${x}_{\text{pad}}$. We then pass ${x}_{\text{pad}}$ through Whisper’s encoder to obtain a feature tensor ${s}_{\text{pad}}\in\mathbb{R}^{1500\times 768}.$ Because only the initial tokens cover the same time span as our acoustic features, we truncate ${s}_{\text{pad}}$ to the first 256 non‐padding time steps: ${s}_{\text{trunc}}\in\mathbb{R}^{256\times 768}.$ To align with the acoustic sequence length, we apply one‐dimensional average pooling over every two consecutive time steps in ${s}_{\text{trunc}}$. This yields the final semantic representation ${O}_{\text{s}}\in\mathbb{R}^{128\times 768}.$ During training, Whisper’s weights remain frozen to preserve high‐level phoneme‐ and word‐level consistency. This design forces the acoustic branch to align its features only when the semantic anchor ${O}_{\text{s}}$, indicates that spoken content is identical.
@@ III-C Semantic‐Acoustic Fusion
The fusion module uses content‐robust semantic embeddings as a stable anchor alongside acoustic features, which capture inherent codec‐related patterns. Inspired by audio-visual synchronization [36, 37], a learnable cross‐attention mechanism [38] then extracts semantic–acoustic mismatch features, which are inspired by audio-visual synchronization. Specifically, three Transformer modules learn these mismatches from different perspectives. Firstly, in the SA Transformer Encoder, semantic queries attend to acoustic keys/values to produce $O_{\text{SA}}$, forcing the semantic stream to highlight codec‐induced distortions. Secondly, in the AS Transformer Encoder, acoustic queries attend to semantic keys/values to produce $O_{\text{AS}}$, downweighting any acoustic components aligned with the anchored content. Finally, these two conditioned maps are passed through a Fusion Transformer that uses $O_{AS}$ as queries and $O_{SA}$ as keys and values to yield $O_{\text{fusion}}$, which is then fed into the AASIST classifier.
@@ III-D Loss Function
where $y$ denotes the input label, $m$ denotes the $margin$, and $L_{MSE_{y}}$ represents the Mean Squared Error between the masked region of the input $e_{\text{coarse}}$ and the corresponding reconstructed region of output $e^{y}_{\text{fine}}$ from $decoder_{y}$. $L_{MSE_{\text{other}}}$ refers to the average MSE between the masked region of $e_{\text{coarse}}$ and corresponding reconstructed masked regions of other outputs $e^{i}_{\text{fine}}$ from $decoder_{i}$, for $i\neq y$. Note that the input $e_{\text{coarse}}$ is masked only during training. During inference, $e_{\text{coarse}}$ is fed directly into the encoder without masking.
@@ IV-A Experimental Setup
Additionally, to assess the effects of unseen speech and silence on performance, we sampled subsets from CoRS: For training, we only use CoRS samples with utterance IDs $\leq 250$, referred to as the CoRS Train (utt. ID $\leq 250$) set. For testing, we split into two subset: a CoRS Test (Seen) set with utterance IDs $\leq 250$ and CoRS Test (Unseen) set $>250$. During training, the model sees only utterances from CoRS Train (utt. IDs $\leq$ 250), while CoRS Test (Unseen) evaluates its ability to generalize to new content. For a detailed discussion, please refer to Section V-A.
We adopt W2V2-AASIST [40]  [^3] https://github.com/TakHemlata/SSL_Anti-spoofing as backbone and train three baseline models on the CoRS Train set under different sampling strategies, referred to as S-VQ, S-AUX, and S-DEC. All input raw waveforms with a length of 82,200 samples at a sampling rate of 16,000 Hz, with RawBoost [41] augmentation. For AudioMAE a patch size of 16 is used, and during training, a masking ratio of 0.4 is applied, divided equally between time and spectral dimensions. No masking is applied during inference. We train models on an NVIDIA RTX 4090 GPU with a batch size of 12 for 40 epochs. The initial learning rate is $5\times 10^{-6}$, and the weight decay is $1\times 10^{-4}$.
@@ V-A The Reason for Model Generalization Failures
Prior work [22] discussed how unseen codec types and generative modeling may affect model generalizability. In this section, we further identify two additional factors, silence and unseen content, that affect model generalization.
@@ V-A1 Silence
One challenge is silence variability: inconsistent pauses at utterance boundaries disrupt codec-signature alignment, causing models to focus on noise in silent segments rather than learning content-independent features. Figure 2(a) shows that CoRS contains silent intervals in both bona fide and resynthesized speech. This occurs because the CoRS samples in CodecFake+ are derived from the untrimmed version of VCTK [42], which includes silence at the beginning and end of each utterance. Since silence is present in both bona fide and spoofed samples of different spoofing algorithms, it may go unnoticed during lab experiments but later leads to degradation when encountering samples with different silence characteristics in the real world. When the model is trained on CoRS with silence at the begining and end of each utterance it can achieve perfect source-tracing F1 scoreson samples with long silences as shown in Figure 2(b). However, the model suffer performance degradation when facing samples with short silences, e.g., from UniAudio [43] and MaskGCT [44]. We further analyzed the correlation between silence and performance by correlating the proportion of silent samples with F1 scores in CoSG ($\rho_{\text{VQ}}=0.62$, $\rho_{\text{AUX}}=0.60$, $\rho_{\text{DEC}}=0.75$). These findings also appear in deepfake detection [45, 46] and highlight the need to model robustly against silence variability.
@@ V-A2 Unseen Content or Speakers
Unseen content and speakers are both notable challenges, which means that linguistic contexts or speaker information not present in the training set introduce phonetic and prosodic patterns outside the model’s learned distribution, thereby reducing performance. Experiments: We conducted a controlled experiment on CoRS set to isolate the effects of silence and variability in speech content on source-tracing performance. During training, the model was trained on the CoRS Train (utt. IDs $\leq$250) set. For evaluation, we used CoRS Test (Seen) set and CoRS Test (Unseen) set to assess the model’s performance on both familiar and unfamiliar speech content. Additionally, we evaluated the impact of silence by comparing model performance on audio with and without silent segments, aiming to determine whether the model had over-fitted to silence.
Results: Table II presents the results of single-task baseline models evaluated on the CoRS Test (Unseen) set along with CoRS Test (Seen) set. Excluding silent segments from the input led to a substantial performance decline, with the single-task model’s F1 score decreasing by approximately 27%. This suggests that the presence of silence has a detrimental effect on the baseline models’ performance. When evaluated on seen speech, the F1 score is approximately 69%, whereas on unseen speech, it decreases to around 67%, indicating that the model’s effectiveness further diminishes when processing unseen speech content. This suggests that the CoSG dataset, which lacks silence and unseen speech components relative to CoRS, poses a challenge for models to generalize to CoSG.
@@ V-B The Impact of Semantic and Acoustic Encoder
Figure 3: Confusion matrices of baseline and SASTNet on CoSG (All) evaluation set, row-normalized by true labels, with predictions on the horizontal axis and true labels on the vertical axis.
@@
Moreover, (Whisper+MAE) integrating Whisper with MAE leads to a slight performance improvement over MAE alone, achieving F1 scores of 41.74 on CoSG (kn.) and 36.94 on CoSG (All). Similarly, incorporating W2V2 with MAE (W2V2+MAE) also enhances performance. Specifically, Pretrained W2V2+M-MAE yields F1 scores of 43.29 on CoSG (kn.) and 36.89 on CoSG (All), while Tuned W2V2+M-MAE achieves 46.18 and 41.91, respectively. These findings indicate that SSL-based features are more robust than traditional mel-spectrograms in handling unseen codecs and generative modeling variations. The superior performance of the Tuned W2V2 could be attributed to its prior knowledge to codec-related data, whereas the Pretrained W2V2 was only pre-trained on bona fide samples. As a result, the Tuned W2V2 is better equipped to capture codec-specific artifacts.
Finally, SASTNet (Pretrained W2V2) achieves F1 scores of 47.28 on CoSG (kn.) and 40.62 on CoSG (All), while SASTNet (Tuned W2V2) achieves 48.13 and 43.47, respectively. These results represent a 24̃F1 improvement over W2V2+MAE, highlighting the effectiveness of fusing refined SSL and Whisper features. This fusion enables the model to better focus on codec-related artifacts with semantic anchors, thereby enhancing performance on CoSG data.
@@ V-C Discussion
To better understanding results, we compared the confusion matrices of SASTNet and the baseline on the CoSG (All) dataset, with the results shown in Fig. 3. Notably, SASTNet demonstrates significantly improved accuracy in detecting unseen real speech across the VQ, AUX, and DEC source tracing tasks. Specifically, on CoSG (All), the accuracy for unseen real speech improves by 2.63$\times$, 2.7$\times$, and 3.36$\times$ compared to the baseline, respectively. Moreover, SASTNet shows improved performance in identifying MVQ, None, Disentanglement, and Frequency-domain codecs, suggesting that it generalizes better across various codec attributes. This can be attributed to its joint modeling of semantic and acoustic features, which allows the model to compare bona fide and codec-resynthesized versions of the same utterance. By leveraging consistent semantic content with contrast acoustic cues, SASTNet is better equipped to capture subtle artifacts left from codecs, thereby enhancing its detection performance.
@@ VI Conclusion
VII Acknowledgements
This work was partially supported by the National Science and Technology Council, Taiwan (Grant no. NSTC 112-2634-F-002-005, Advanced Technologies for Designing Trustable AI Services). We also thank the National Center for High-performance Computing (NCHC) of National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources.

### v3

@@ I Introduction
Deepfake detection determines whether the given speech is a bona fide speech or a deepfake speech. Numerous challenges and datasets [1, 2] have been introduced, such as ASVspoof [3, 4, 5, 6, 7] and ADD [8, 9]. Recently, attention has shifted from merely detecting deepfake speech to tracing its source. Source tracing offers important insights into the origin or generation algorithm of a deepfake speech, which is especially valuable in out‐of‐domain detection scenarios. The simplest approach to source tracing [10] involves mapping deepfake utterances back to the specific attack model ID that generated them [9]. To enhance generalization, the field has shifted toward attribute‐based source tracing [11, 12, 13, 14]. Rather than treating each deepfake system as a unique category, this method groups systems according to shared attributes, reflecting the fact that many deepfake pipelines reuse common components. By focusing on these attributes, models can more effectively trace spoofing algorithms that were unseen during training but are built from known elements during training [14]. Drawing on well‐defined properties from traditional text-to-speech (TTS) and voice conversion (VC) architectures, researchers have concentrated on classifying attributes such as input types [14], acoustic models [12, 14], speaker representations [13], and vocoders [13, 11, 12, 14], etc.
However, these prior studies target source tracing in traditional TTS/VC systems, and only initial efforts have explored codec‐based speech generation (CoSG) [15] models. Xie et al. [16] treat source tracing as a closed‐set classification problem—identifying known CoSG system IDs and using a binary out‐of‐distribution detector, but this approach fails when multiple unseen generators are encountered. Although there has been a series of recent works on source tracing [17, 18, 19, 20], few focus on codec-based deepfake speech [21, 22]. More recently, Chen et al. [23] utilize the organized CodecFake+[22] taxonomy, which covers vector quantization schemes, auxiliary objectives, and decoder types, to formulate three source‐tracing tasks grounded in codec attributes. They [23] also show that a model trained on codec‐resynthesized data suffers from severe generalization issues when evaluated on CoSG‐generated speech, driven by unseen codec types.
@@ II Background
Neural audio codecs [24, 25, 15, 26, 27] leave “fingerprints” across three axes: vector quantization, auxiliary objectives, and decoder type. Together, these axes span most codec designs and enable robust tracing of codec-based deepfakes. Building on the taxonomy, Chen et. al. [23] define three multi-class classification tasks for CodecFake source tracing:
  • Vector Quantization Classification (VQ Task): Four-class classification: classifies a waveform as real or generated by a model with codec using multi-codebook (Mvq), single-codebook (Svq), or scalar quantization (Scq).
  • Auxiliary Objective Classification (AUX Task): Four-class classification: classifies a waveform as either real or generated by a model with a codec using semantic distillation (Sem), disentanglement objectives (Disent), or no auxiliary objective (None).
  • Decoder Type Classification (DEC Task): Three-class classification: classifies an input utterance as real or generated by a model with a codec employing a time- or frequency-domain decoder (denoted Time and Freq).
Note that we focus solely on the source-tracing tasks above, not on traditional anti-spoofing.
@@ III-A Semantic Encoder
The Semantic Encoder leverages a frozen, pre‐trained Whisper [30] model to generate content‐robust embeddings that serve as a stable anchor for source tracing. By verifying that two utterances share identical linguistic content, these embeddings ensure the network ignores natural acoustic variations caused by different words or pronunciations and focuses only on codec‐induced distortions. Concretely, let x{x} denote the raw input waveform. We first zero‐pad x{x} to a fixed duration of 30 seconds, yielding ${x}_{\text{pad}}$. We then pass ${x}_{\text{pad}}$ through Whisper’s encoder to obtain a feature tensor ${s}_{\text{pad}}\in\mathbb{R}^{1500\times 768}.$ Because only the initial tokens cover the same time span as our acoustic features, we truncate ${s}_{\text{pad}}$ to the first 256 non‐padding time steps: ${s}_{\text{trunc}}\in\mathbb{R}^{256\times 768}.$ To align with the acoustic sequence length, we apply one‐dimensional average pooling over every two consecutive time steps in ${s}_{\text{trunc}}$. This yields the final semantic representation ${O}_{\text{s}}\in\mathbb{R}^{128\times 768}.$ During training, Whisper’s weights remain frozen to preserve high‐level phoneme‐ and word‐level consistency. This design forces the acoustic branch to align its features only when the semantic anchor ${O}_{\text{s}}$ indicates that spoken content is identical.
@@ III-C Semantic‐Acoustic Fusion
The fusion module uses content‐robust semantic embeddings as an anchor with acoustic features, which capture inherent codec‐related patterns. Inspired by audio-visual learning [37, 38, 39], a learnable cross‐attention mechanism [40] then extracts semantic–acoustic mismatch features, which are inspired by audio-visual synchronization. Specifically, three modules learn these mismatches from different aspects. Firstly, in the SA Transformer Encoder, semantic queries attend to acoustic keys/values to produce $O_{\text{SA}}$, forcing the semantic stream to highlight codec‐induced distortions. Secondly, in the AS Transformer Encoder, acoustic queries attend to semantic keys/values to produce $O_{\text{AS}}$, down-weighting any acoustic components aligned with the anchored content. Finally, these two conditioned maps are passed through a Fusion Transformer that uses $O_{\text{AS}}$ as queries and $O_{\text{SA}}$ as keys and values to yield $O_{\text{fusion}}$, which is then fed into the AASIST classifier.
@@ III-D Loss Function
where $y$ denotes the input label, $m$ denotes the marginmargin, and $L_{MSE_{y}}$ represents the Mean Squared Error between the masked region of the input $e_{\text{coarse}}$ and the corresponding reconstructed region of output $e^{y}_{\text{fine}}$ from $decoder_{y}$. $L_{MSE_{\text{other}}}$ refers to the average MSE between the masked region of $e_{\text{coarse}}$ and corresponding reconstructed masked regions of other outputs $e^{i}_{\text{fine}}$ from $decoder_{i}$, for $i\neq y$. Note that the input $e_{\text{coarse}}$ is masked only during training. During inference, $e_{\text{coarse}}$ is fed directly into the encoder without masking.
@@ IV-A Experimental Setup
Additionally, to assess the effects of unseen speech and silence on performance, we sampled subsets from CoRS: For training, we only use CoRS samples with utterance IDs $\leq 250$, referred to as the CoRS Train (utt. ID $\leq 250$) set. For testing, we split into two subset: a CoRS Test (Seen) set with utterance IDs $\leq 250$ and CoRS Test (Unseen) set >250>250. During training, the model sees only utterances from CoRS Train (utt. IDs $\leq$ 250), while CoRS Test (Unseen) evaluates its ability to generalize to new content. For a detailed discussion, please refer to Section V-A.
We adopt W2V2-AASIST [42]  [^2] https://github.com/TakHemlata/SSL_Anti-spoofing as backbone and train three baseline models on the CoRS Train set under different sampling strategies, referred to as S-VQ, S-AUX, and S-DEC. All input waveforms with a length of 82,200 samples at a sampling rate of 16,000 Hz, with RawBoost [43] augmentation. For AudioMAE, a patch size of 16 is used, and during training, a masking ratio of 0.4 is applied, divided equally between time and spectral dimensions. No masking is applied during inference. We train models on an NVIDIA RTX 4090 GPU with a batch size of 12 for 40 epochs. We use the Cosine Annealing learning rate (LR) scheduler with the initial learning rate is $5\times 10^{-6}$, and the weight decay is $1\times 10^{-4}$.
@@ V-A The Reason for Model Generalization Failures
Prior work [23] discussed how unseen codec types and generative modeling may affect model generalization. In this section, we further identify two additional factors, silence and unseen content, that affect model generalization.
@@ V-A1 Silence
One challenge is silence variability: inconsistent pauses at utterance boundaries disrupt codec-signature alignment, causing models to focus on noise in silent segments rather than learning content-independent features. Figure 2(a) shows that CoRS contains silent intervals in both bona fide and resynthesized speech. This occurs because the CoRS samples in CodecFake+ are derived from the untrimmed version of VCTK [44], which includes silence at the beginning and end of each utterance. Since silence is present in both bona fide and spoofed samples of different spoofing algorithms, it may go unnoticed during lab experiments, but later leads to degradation when encountering samples with different silence characteristics in the real world. When the model is trained on CoRS with silence at the beginning and end of each utterance, it can achieve perfect source-tracing F1 scores on samples with long silences as shown in Figure 2(b). However, the model suffers performance degradation when facing samples with short silences, e.g., from UniAudio [45] and MaskGCT [46]. We further analyzed the correlation between silence and performance by correlating the proportion of silent samples with F1 scores in CoSG ($\rho_{\text{VQ}}=0.62$, $\rho_{\text{AUX}}=0.60$, $\rho_{\text{DEC}}=0.75$). These findings also appear in deepfake detection [47, 48] and highlight the need to model robustly against silence variability.
@@ V-A2 Unseen Content or Speakers
Unseen content and speakers are both notable challenges, which means that linguistic contexts or speaker information not present in the training set introduce phonetic and prosodic patterns outside the model’s learned distribution, thereby reducing performance. Experiments: We conducted a controlled experiment on CoRS set to isolate the effects of silence and variability in speech content on source-tracing performance. During training, the model was trained on the CoRS Train (utt. IDs $\leq$250) set. For evaluation, we used CoRS Test (Seen) set and CoRS Test (Unseen) set to assess the performance of the model in both familiar and unfamiliar speech content. Additionally, we evaluated the impact of silence by comparing model performance on audio with and without silent segments, aiming to determine whether the model had overfitted to silence.
Results: Table II presents the results of single-task baseline models evaluated on the CoRS Test (Unseen) set, along with the CoRS Test (Seen) set. Excluding silent segments from the input led to a substantial performance decline, with the single-task model’s F1 score decreasing by approximately 27%. This suggests that the presence of silence has a detrimental effect on the baseline models’ performance. When evaluated on seen speech, the F1 score is approximately 69%, whereas on unseen speech, it decreases to around 67%, indicating that the model’s effectiveness further diminishes when processing unseen speech content. This suggests that the CoSG dataset, which lacks silence and unseen speech components relative to CoRS, poses a challenge for models to generalize to CoSG.
@@ V-B The Impact of Semantic and Acoustic Encoder
Figure 3: Confusion matrices of baseline and SASTNet on CoSG (All) evaluation set, row-normalized by true labels, with predictions on the horizontal axis and true labels on the vertical axis. More details about labels in Section II
@@
Moreover, (Whisper+MAE) integrating Whisper with MAE leads to a slight performance improvement over MAE alone, achieving F1 scores of 41.74 on CoSG (kn.) and 36.94 on CoSG (All). Similarly, incorporating W2V2 with MAE (W2V2+MAE) also enhances performance. Specifically, Pretrained W2V2+M-MAE yields F1 scores of 43.29 on CoSG (kn.) and 36.89 on CoSG (All), while Tuned W2V2+M-MAE achieves 46.18 and 41.91, respectively. These findings indicate that SSL-based features are more robust than traditional mel-spectrograms in handling unseen codecs and generative modeling variations. The superior performance of the Tuned W2V2 could be attributed to its prior knowledge of codec-related data, whereas the Pretrained W2V2 was only pre-trained on bona fide samples. As a result, the Tuned W2V2 is better equipped to capture codec-specific artifacts.
Finally, SASTNet (Pretrained W2V2) achieves F1 scores of 47.28 on CoSG (kn.) and 40.62 on CoSG (All), while SASTNet (Tuned W2V2) achieves 48.13 and 43.47, respectively. These results represent a 24̃% F1 improvement over W2V2+MAE, highlighting the effectiveness of fusing refined SSL and Whisper features. This fusion enables the model to better focus on codec-related artifacts with semantic anchors, thereby enhancing performance on CoSG data.
@@ V-C Discussion
[…]
@@ VI Conclusion
VII Acknowledgments
We thank National Center for High-performance Computing (NCHC) of National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources.